OpenStudy (anonymous):

(x^6-x^5)(x^6+x^5) multiply

OpenStudy (anonymous):

x^12 - x^10 = x^10(x^2 - 1)

OpenStudy (anonymous):

x^12 - x^10 = x^10(x^2 - 1) = x^10(x - 1)(x + 1)

OpenStudy (anonymous):

@unklerhaukus your answer is not right.

OpenStudy (unklerhaukus):

@GT i see now what i did wrong \[(x^6-x^5)(x^6+x^5)=x^5(x-1)x^5(x-1)\]\[=x^{10}(x-1)(x+1)=x^{10}(x^2-1)\]\[=x^{12}-x^{10}\]